how can I put a real world problemin to the equation 70 + 5x = 40 + 10x

Answer by addingup(3677)

We make skateboards. Our client supplies the wood, wheels and other materials.
:
The cost of making skateboards in an automated machine that costs $70 and takes $5 of labor to operate is the same as making skateboards in a cheaper machine that costs $40 but needs more labor to operate, it needs $10 of labor:
70 + 5x = 40 + 10x, where x is the number of units produced.
:
Solve the equations: 70-40 = 10x-5x = 30 = 5x; x = 6 At 6 skateboards our production cost using either machine is the same. Under 6, using the $40 machine is cheaper. But over 6 using the $70 machine is cheaper. How many units do you forecast our customer will want?